6X+8y=10
26x+48y=2
sir/mam, how to find the value of x and y....pls rply with full explaination..what isthe basic thing need to do for solving linear equation..pls rply.....am very poor in mathematics..

Answer by kibauu(44)

(1)6X+8y=10
(2)26x+48y=2
the basic thing to solve linear is making x variabel or y variabel gone.
In eq.1, you have 8y, and in eq.2 you have 48y. To solve it, multiply eq.1 with 6, so it will be 36x+48y=60. Look, y has the same variabel. All you have to do is subtracting it.
36x+48y=60
26x+48y=2
--------- -
10x=58
then, it's easy to find x
x=5,8
Practice makes perfect. Dont forget to read your textbook or google it first.
And yeah sorry for my english, i'm not a native English spearker :)
Cheers